OpenStudy (kkutie7):

I've forgotten how to set this problem up:

OpenStudy (kkutie7):

If \[\int\limits_{1}^{3}f(w)dw=7\] find the value of \[\int\limits_{1}^{2}f(5-2x)dx\]

OpenStudy (anonymous):

Substitute \[5-2x=w\]

random231 (random231):

by f(omega) it means a periodic function right?

OpenStudy (kkutie7):

uh.... I don't think so

random231 (random231):

nishant is right!

OpenStudy (kkutie7):

@Nishant_Garg that doesn't work

OpenStudy (kkutie7):

Would work if the limits were the same

random231 (random231):

that does! 5-2x=w then -2dx=dw you need to change the limits then

OpenStudy (anonymous):

\[5-2x=w\]\[dx=-\frac{1}{2}dw\] Limits become 3 to 1